OpenStudy (anonymous):

3 + 6 = 9 Choose the related addition sentence that shows the commutative property of addition.

OpenStudy (anonymous):

6+3=9, just switch the numbers.

OpenStudy (anonymous):

Commutative Law says that If a + b = c then b + a = c a + b = b+c So 3 + 6 = 9 6 + 3 =9

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Selena orders CDs for a music shop. She just ordered 8 CDs from each of 11 pop artists and 5 CDs from each of 8 jazz artists. How many CDs did she order altogether?